OpenStudy (courtstacks98):

How do you prove the following trigonometric identity? sec(theta)/tan(theta)=csc(theta)

OpenStudy (jtug6):

Had to look this one up im a bit rusty on le trig but you should first try to get everything in terms of sin/cos and perhaps get rid of that nasty fraction :P

OpenStudy (raegan360brown):

sec= 1/cos tan= sin/cos csc= 1/sin now write this out and simplify

OpenStudy (jtug6):

Basically that.

OpenStudy (courtstacks98):

Thank you both!

OpenStudy (jtug6):

Np :)